a man is 67. His son is 29. In how many years will his son be exactly half his father’s age?
Answer by jorel1380(3719)

29+y=1/2(67+y)
29+y=67/2+y/2
58+2y=67+y
y=67-58=9
The son will be half the father's age in 9 years.
38=76/2..